OpenStudy (alina123):

The sum of R,S,T is 40. The ratio of R to S is 1:3. The ratio of S to T is 3:4. What is S? ( This was a question on the SHSAT i just took yesterday) @Loser66 @Destinymasha @iambatman

OpenStudy (loser66):

R+ S + T= 40 \(\dfrac{R}{S}= \dfrac{1}{3}\rightarrow R=\dfrac{1}{3}S\) \(\dfrac{S}{T}= \dfrac{3}{4}\rightarrow T=\dfrac{4}{3}S\) just plug all into the first one to solve for S